# The neutrino luminosity and energy spectrum of nova outburst  [^1] Released on March, 1st, 2021

### v1

@@ 3 Result
Using MESA, we simulate the structure and evolution of 7 nova models including different initial CO WD masses from 0.6 to 1.1 M with a mass interval of 0.1 M, and 1.1 M ONeMg WD. Following Farag et al. (2020), we utilize the neutrino H-R diagram to illustrate the neutrino luminosities throughout the complete life cycles of nova outburst, and discuss how $M_{\mathrm{WD}}$, $\dot{M}$, $T_{\mathrm{C}}$, and mixing affect low-energy nuclear neutrino luminosity of nova.
3.1 neutrino luminosity of nova outburst
@@
Figure 1 shows the nova with 0.9 M CO WD and 1.1 M ONeMg WD singly periodic evolutionary tracks in both the H-R diagram and the neutrino H-R diagram. Several key moments during the nova eruption have been marked in the H-R diagram and neutrino H-R diagram. The $a\to b\to c\to d$ is the explosion curve, the $d\to e\to a$ is the cooling curve. The a $\to$ c clearly illustrates that when the TNR occurs, due to the rapid rise of temperature the intense nuclear reaction produce a large number of neutrinos, and the cross-section of neutrinos interacting with matter is extremely low compared to photons, the low-energy nuclear and thermal neutrino luminosity rises rapidly in a straight line while photon luminosity remains almost constant. It is why in the neutrino H-R diagram, there are two points labeled $a$. When the low-energy nuclear and thermal neutrino luminosity reaches its maximum, it also means that the nuclear reaction rate on the surface of the WD reaches its maximum, the photon luminosity begins to rise, low-energy nuclear and thermal neutrino luminosity would remain constant. The c $\to$ d indicates that when part of the photon’s energy is injected into the shell, the shell begins to expand, the surface temperature of the WD decreases. The particle number density in the nuclear reaction zone decreases, and the nuclear reaction rate begins to decrease, so the low-energy nuclear and thermal neutrino luminosity begins to decrease. The d $\to$ e indicates that photosphere of the WD begins to contract, the effective temperature rises, the nuclear reaction rate remains constant, as does the photon luminosity and low-energy nuclear and thermal neutrino luminosity. Finally, the nova outburst is over. The cooling process of the WD commences as the nuclear reaction concludes, leading to a gradual decline in photon and low-energy nuclear and thermal neutrino luminosity.
@@
Figure 1: The track in the H-R diagram and neutrino H-R diagram for the whole nova multicycle evolution. The left is 0.9 M CO nova model, the right is 1.1 M ONe nova model. Both have center temperature $T_{\mathrm{C}}=3\times 10^{7}\ \rm K$, accrete $25\%$ WD material and $75\%$ solar material at a rate $1\times 10^{-9}\ \mathrm{M}_{\odot}\rm\ yr^{-1}$.
@@
The production of neutrinos can be achieved through both nuclear reactions and thermal processes, as discussed in section 2.1. The left panel of Figure 2 illustrates the ratio of the low-energy nuclear reaction neutrino luminosity to the low-energy thermal process neutrino luminosity, traced along the light curve of the nova eruption. The right panel of Figure 2 shows the ratio of the low-energy nuclear and thermal neutrino luminosity to the photon luminosity, traced along the light curve of the nova eruption. During the accretion phase, the temperature of the accretion layer is less than the Fermi temperature $(7\times 10^{7}\rm\ K)$, the nuclear reactions are weak, neutrinos are mainly produced by thermal processes, the accretion layer of the WD radiates energy mainly through photons. However, when the TNR occurs, the neutrinos produced by nuclear reactions will far exceed those produced by thermal processes. Thus, neutrinos loss becomes the main channel of energy loss at the TNR. In addition, low-energy nuclear and thermal neutrino luminosity far exceeds photon luminosity when the nova outbursts. For example, as our $M_{\mathrm{CO}}=0.9\ \rm M_{\odot}$ nova model, when it outbursts, the peak photon luminosity is $5.5\times 10^{4}\ \rm L_{\odot}$, the peak low-energy nuclear neutrino luminosity is up to $3.6\times 10^{6}\ \rm L_{\gamma,\odot}$.
@@
Figure 2: The low-energy neutrino luminosity produced by the nuclear reaction and thermal processes (the left panel). The colorful solid lines are the ratio of the low-energy nuclear reaction neutrino luminosity to the low-energy thermal neutrino luminosity plotted along the light curve of the nova eruption. Gray curves indicate where nuclear reaction neutrinos dominate, green curves where thermal neutrinos dominate, and blue and red curves where the reaction and thermal neutrino luminosities are within a factor of 10. The right panel shows the ratio of the low-energy nuclear and thermal neutrino luminosity to the photon luminosity during the whole nova multicycle evolution. The colorful solid lines give the ratio of the low-energy nuclear and thermal neutrino luminosity to the photon luminosity in a light curve of the nova eruption. The nova model has a 0.9 $\rm M_{\odot}$ CO WD, center temperature $T_{\mathrm{C}}=3\times 10^{7}\ \rm K$, accrete $25\%$ WD material and $75\%$ solar material at a rate $1\times 10^{-9}\ \mathrm{M}_{\odot}\rm\ yr^{-1}$.
@@
Figure 3 shows the the energy yields of the four thermal neutrino processes, and the nuclear reaction rates of the nine neutrinos producing nuclear reactions in the pp chain and the CNO cycle. The left panel shows that when the pre-TNR accreted mass is determined, for the nuclear process, the proton-proton chain plays a crucial role during the primary accretion phase, especially $pp$ reaction: $\rm p+p\to d+e^{+}+\nu$ (Starrfield et al., 2016) and $\rm{}^{7}Be$ electron capture :$\rm{}^{7}Be(e^{-},\nu)^{7}Li$. As accretion progresses, when the temperature of layers is about $7\times 10^{7}\ \rm K$, the degeneracy is no longer important and the layers begin to expand, the temperature begins to rise rapidly, nuclear reaction neutrinos are dominated by CNO cycle, and the TNR occurs. Neutrinos are produced mainly by four nuclear reactions: $\rm{}^{13}N(e^{+},\nu_{e})^{13}C$, $\rm{}^{15}O(e^{+},\nu_{e})^{15}N$, $\rm{}^{17}F(e^{+},\nu_{e})^{17}O$, $\rm{}^{18}F(e^{+},\nu_{e})^{18}O$. For thermal process, during the accretion phase, the WD has a high density in both its interior and layers, the neutrino production is dominated by plasma decay. Once the TNR occurs, neutrino production will be dominated by photon-neutrino because the temperature of layers is in the temperature range $\rm(T<4\times 10^{8}\ K)$ that it dominates, as shown in the right panel.
@@
Figure 3: The nuclear reactions rates of the nine neutrinos producing nuclear reactions in the pp chain and the CNO cycle during the whole nova multicycle evolution (the left panel). The right panel shows the energy yield of four thermal processes during the whole nova multicycle evolution. The blue line is plasma decay, yellow line is bremsstrahlung, green line is photon-neutrino, cyan line is electro pair annihilation. The nova model is similar to Figure 2.
@@
3.2 Neutrino luminosity of novae with different $M_{\mathrm{WD}}$, accretion rate $\dot{M}$, $T_{\mathrm{C}}$, and mixing
@@
The $M_{\mathrm{WD}}$ of nova can affect temperatures in the nuclear burning region. The left panel of Figure 4 shows the neutrino luminosity curves of the range of CO $M_{\mathrm{WD}}$ is $0.6\ \rm M_{\odot}$ to $1.1\ \rm M_{\odot}$. The right panel shows the average peak low-energy nuclear neutrino luminosity of 10 outbursts. It is clearly illustrates that the peak luminosity of neutrinos increases as the mass of the WD increases. The larger the mass and smaller the radius of a WD, the greater the gravitational acceleration. This makes it harder for hydrogen to penetrate the degenerate WD core. That means a more degenerate hydrogen-rich envelope, higher temperatures, and, of course, a more violent explosion. As discussed above, nuclear neutrino is mainly produced by CNO cycle, CNO cycle is highly sensitive to template ($\varepsilon_{\rm CNO}\propto T^{17}$). So the more massive the $M_{\mathrm{WD}}$, the higher the low-energy nuclear neutrino luminosity during the TNR.
@@
Figure 4: Neutrino luminosity curves with different CO nova mass models. Left: The neutrino luminosity of the burst sequence during 0-400000 year with different CO nova mass models, i.e., 0.6 $\rm M_{\odot}$ (cyan), 0.7 $\rm M_{\odot}$ (grey), 0.8 $\rm M_{\odot}$ (red), 0.9 $\rm M_{\odot}$ (green), 1.0 $\rm M_{\odot}$ (blue), and 1.1 $\rm M_{\odot}$ (yellow). Right: Average light curves. Both nova models have center temperature $T_{\mathrm{C}}=3\times 10^{7}\ \rm K$, accrete $25\%$ WD material and $75\%$ solar material at a rate $\dot{M}=1\times 10^{-9}\ \mathrm{M}_{\odot}\rm\ yr^{-1}$.
@@
The left panel of Figure 5 shows the neutrino luminosity curves of nova models with different $\dot{M}$. The right panel shows the average peak low-energy nuclear neutrino luminosity of 10 outbursts. It is clearly show that the higher the accretion rate $\dot{M}$, the smaller the outburst interval $\Delta t_{\mathrm{rec}}$ and $L_{\nu}$. TNR is produced when the hydrogen-rich material is heated to a high enough temperature $\rm(7\times 10^{7}\ K)$ for the hydrogen to burn through the CNO cycle. The high accretion rate provides heat at a high rate (Prialnik et al., 1982), thereby inducing an acceleration in nova eruptions. However, owing to the reduced mass of accretion, the lower TNR peak temperature eventually gives rise to a diminished low-energy nuclear neutrino luminosity at higher accretion rates compared to lower ones. When the accretion rate is low, the accretion layer temperature rises slowly and the outburst interval is larger, which means that the accretion layer accrets more material, the outburst intensity is higher, and the low-energy nuclear neutrino luminosity produced during the outburst is higher.
@@
Figure 5: Neutrino luminosity curves with different $\dot{M}$. Left: The neutrino luminosity of the burst sequence during 0-300000 year with different $\dot{M}$, i.e., $1\times 10^{-9}\ \mathrm{M}_{\odot}\rm\ yr^{-1}$ (cyan), $2\times 10^{-9}\ \mathrm{M}_{\odot}\rm\ yr^{-1}$ (grey), $3\times 10^{-9}\ \mathrm{M}_{\odot}\rm\ yr^{-1}$ (red), $4\times 10^{-9}\ \mathrm{M}_{\odot}\rm\ yr^{-1}$ (green). Right: Average light curves. Both nova models have center temperature $T_{\mathrm{C}}=3\times 10^{7}\ \rm K$, accrete $25\%$ WD material and $75\%$ solar material at a 0.9 $\rm M_{\odot}$ CO WD.
@@
The left panel of Figure 6 displays the neutrino luminosity curves of a hot model $(T_{\mathrm{C}}=3\times 10^{7})$ and cold models $(T_{\mathrm{C}}=2\times 10^{7}\ \mathrm{K}\ \mathrm{and}\ T_{\mathrm{C}}=1\times 10^{7}\ \rm K)$ and with $0.9M_{\odot}$ WD. The right panel shows the average peak low-energy nuclear neutrino luminosity of 10 outbursts. When the initial core temperature of the nova model WD is lower, most of the gravitational energy released by the accretion compressed material will be directed to the core, which will delay the time to reach the TNR. Of course, it means the envelope will accrete more H-rich materials, the outbursts are more intense and the low-energy nuclear neutrino luminosity is higher. However, due to the lower temperature of the core and accreted layer, the nuclear reaction rate during accretion is also lower, resulting in a low neutrino luminosity during the accrete phase. In the hot WD $(T_{\mathrm{C}}=3\times 10^{7}\ \rm K)$, convection zones will appear in the outer core layer. The accreted hydrogen rapidly mixes in the convective region and continues to diffuse to the core, leading to an early outburst (Schwartzman et al., 1994). However, because the accretion mass is less and the temperature is lower, the low-energy nuclear neutrino luminosity produced by the outburst is lower.
@@
Figure 6: Neutrino luminosity curves with different $T_{\mathrm{C}}$. Left: The neutrino luminosity of the burst sequence during 0-50000 year with different $T_{\mathrm{C}}$, i.e., $1\times 10^{7}\rm\ K$ (red), $2\times 10^{7}\rm\ K$ (grey), $3\times 10^{7}\rm\ K$ (cyan). Right: Average light curves. Both nova models have accrete rate $1\times 10^{-8}\ \mathrm{M}_{\odot}\rm\ yr^{-1}$, accrete $25\%$ WD material and $75\%$ solar material at a 0.9 $\rm M_{\odot}$ CO WD.
@@
The left panel of Figure 7 shows the neutrino luminosity curves of nova models when mixing is considered or not. The right panel shows the average peak low-energy nuclear neutrino luminosity of 10 outbursts. If the mixing of accretion material and core CO-enriched material is considered during the accretion stage, it means an increase in the metallicity of the accretion material. On the one hand, increasing the metallicity of the accreting material results in an increase in the opacity. The increased opacity not only allows more heat from the compression of accretion material to be trapped in the accretion layer, but also more heat from nuclear burning to be trapped in the nuclear reaction zone so that the temperature in the nuclear burning region increases faster and the shorter the time to reach TNR than ignore mixing nova model. Therefore, the total amount of accreted and ejected mass is less (Starrfield et al., 1998). On the other hand, mixing of CO-enriched materials can also outburst earlier because it acts as a catalyst for the CNO cycle (H burning). Models that do not take into account mixing require more accretion matter because the temperature rises more slowly in the nuclear burning region. Thus, compared to nova models that consider mixing, the nova model that does not consider mixing has a denser accretion layer, more intense outbursts, and higher low-energy nuclear neutrino luminosity.
@@
Figure 7: Neutrino luminosity curves with different mixing. Left: The neutrino luminosity of the burst sequence during 0-15000 year when accrete $25\%$ WD material and $75\%$ solar material and not consider any mixing. Right: Average light curves. Both nova models have center temperature $T_{\mathrm{C}}=3\times 10^{7}\ \rm K$, 0.8 $\rm M_{\odot}$ CO WD at a rate $\dot{M}=1\times 10^{-8}\ \mathrm{M}_{\odot}\rm\ yr^{-1}$.
3.3 neutrino luminosity of T CrB nova
Considering $M_{\rm WD}$ of T CrB is 1.1-1.3 $\rm M_{\odot}$ and the time span between outbursts is approximately 80 years, the average accretion rate $\dot{M}_{a}\approx 2\times 10^{-8}\ \mathrm{M}_{\odot}\rm yr^{-1}$ (Zamanov et al., 2023), so we use $M_{\mathrm{WD}}=1.3\ \rm M_{\odot}$, $\dot{M}=1.72\times 10^{-8}\ \mathrm{M}_{\odot}\rm\ yr^{-1}$ as our T CrB nova model. The left panel of Figure 8 simulates the low-energy nuclear neutrino luminosity of the first two bursts of the T CrB and the next one shortly after, with peak low-energy nuclear neutrino luminosity of $1.06\times 10^{8}\ L_{\nu,\odot}$, $2.05\times 10^{8}\ L_{\nu,\odot}$, and $2.70\times 10^{8}\ L_{\nu,\odot}$, respectively. In addition, According to the definition of outburst duration in section 2, it can be predicted that the next outburst of T CrB will last about 88 days, as shown in the right panel of Figure 8.

### v2

@@ III Result
Using mesa, we simulate the structure and evolution of seven nova models including different initial CO WD masses from 0.6 to 1.1 M with a mass interval of 0.1 M, and 1.1 M ONeMg WD. Following Farag et al. [67], we utilize the neutrino H-R diagram to illustrate the neutrino luminosities throughout the complete life cycles of nova outburst and discuss how $M_{\mathrm{WD}}$, $\dot{M}$, $T_{\mathrm{C}}$, and mixing affect low-energy nuclear neutrino luminosity of the nova.
III.1 Neutrino luminosity of nova outburst
@@
Figure 1 shows the nova with 0.9 M CO WD and 1.1 M ONeMg WD singly periodic evolutionary tracks in both the H-R diagram and the neutrino H-R diagram. Several key moments during the nova eruption have been marked in the H-R diagram and neutrino H-R diagram. The $a\to b\to c\to d$ is the explosion curve, and the $d\to e\to a$ is the cooling curve. The $a\to c$ clearly illustrates that, when the TNR occurs, due to the rapid rise of temperature, the intense nuclear reaction produces a large number of neutrinos, the cross section of neutrinos interacting with the matter is extremely low compared to photons, and the low-energy nuclear and thermal neutrino luminosity rises rapidly in a straight line, while photon luminosity remains almost constant. This is why, in the neutrino H-R diagram, there are two points labeled $a$. When the low-energy nuclear and thermal neutrino luminosity reaches its maximum, it also means that the nuclear reaction rate on the surface of the WD reaches its maximum, the photon luminosity begins to rise, and low-energy nuclear and thermal neutrino luminosity would remain constant. The $c\to d$ indicates that when part of the photon’s energy is injected into the shell, the shell begins to expand, and the surface temperature of the WD decreases. The particle number density in the nuclear reaction zone decreases, and the nuclear reaction rate begins to decrease, so the low-energy nuclear and thermal neutrino luminosity begins to decrease. The $d\to e$ indicates that photosphere of the WD begins to contract, the effective temperature rises, and the nuclear reaction rate remains constant, as does the photon luminosity and low-energy nuclear and thermal neutrino luminosity. Finally, the nova outburst is over. The cooling process of the WD commences as the nuclear reaction concludes, leading to a gradual decline in photon and low-energy nuclear and thermal neutrino luminosity.
@@
Figure 1: The track in the H-R diagram and neutrino H-R diagram for the whole nova multicycle evolution. The left is 0.9 M CO nova model, the right is 1.1 M ONe nova model. Both have center temperature $T_{\mathrm{C}}=3\times 10^{7}\ \rm K$ and accrete $25\%$ WD material and $75\%$ solar material at a rate $1\times 10^{-9}\ \mathrm{M}_{\odot}\rm\ yr^{-1}$.
@@
The production of neutrinos can be achieved through both nuclear reactions and thermal processes, as discussed in Sec. II.1. The left panel of Fig. 2 illustrates the ratio of the low-energy nuclear reaction neutrino luminosity to the low-energy thermal process neutrino luminosity, traced along the light curve of the nova eruption. The right panel of Fig. 2 shows the ratio of the low-energy nuclear and thermal neutrino luminosity to the photon luminosity, traced along the light curve of the nova eruption. During the accretion phase, the temperature of the accretion layer is less than the Fermi temperature $(7\times 10^{7}\rm\ K)$, the nuclear reactions are weak, neutrinos are mainly produced by thermal processes, and the accretion layer of the WD radiates energy mainly through photons. However, when the TNR occurs, the neutrinos produced by nuclear reactions will far exceed those produced by thermal processes. Thus, neutrinos loss becomes the main channel of energy loss at the TNR. In addition, low-energy nuclear and thermal neutrino luminosity far exceeds photon luminosity when the nova outbursts. For example, as our $M_{\mathrm{CO}}=0.9\ \rm M_{\odot}$ nova model, when it outbursts, the peak photon luminosity is $5.5\times 10^{4}\ \rm L_{\odot}$, and the peak low-energy nuclear neutrino luminosity is up to $3.6\times 10^{6}\ \rm L_{\gamma,\odot}$.
@@
Figure 2: The low-energy neutrino luminosity produced by the nuclear reaction and thermal processes (left). The colorful solid lines are the ratio of the low-energy nuclear reaction neutrino luminosity to the low-energy thermal neutrino luminosity plotted along the light curve of the nova eruption. Gray curves indicate where nuclear reaction neutrinos dominate, green curves where thermal neutrinos dominate, and blue and red curves where the reaction and thermal neutrino luminosities are within a factor of 10. Right: the ratio of the low-energy nuclear and thermal neutrino luminosity to the photon luminosity during the whole nova multicycle evolution. The colorful solid lines give the ratio of the low-energy nuclear and thermal neutrino luminosity to the photon luminosity in a light curve of the nova eruption. The nova model has a 0.9 $\rm M_{\odot}$ CO WD, center temperature $T_{\mathrm{C}}=3\times 10^{7}\ \rm K$, and accretes $25\%$ WD material and $75\%$ solar material at a rate $1\times 10^{-9}\ \mathrm{M}_{\odot}\rm\ yr^{-1}$.
@@
Figure 3 shows the energy yields of the four thermal neutrino processes and the nuclear reaction rates of the nine neutrinos producing nuclear reactions in the pp chain and the CNO cycle. The left panel shows that when the pre-TNR accreted mass is determined, for the nuclear process, the proton-proton chain plays a crucial role during the primary accretion phase, especially the $pp$ reaction: $\rm p+p\to d+e^{+}+\nu$ [63] and $\rm{}^{7}Be$ electron capture :$\rm{}^{7}Be(e^{-},\nu)^{7}Li$. As accretion progresses, when the temperature of layers is about $7\times 10^{7}\ \rm K$, the degeneracy is no longer important and the layers begin to expand, the temperature begins to rise rapidly, nuclear reaction neutrinos are dominated by the CNO cycle, and the TNR occurs. Neutrinos are produced mainly by four nuclear reactions: $\rm{}^{13}N(e^{+},\nu_{e})^{13}C$, $\rm{}^{15}O(e^{+},\nu_{e})^{15}N$, $\rm{}^{17}F(e^{+},\nu_{e})^{17}O$, and $\rm{}^{18}F(e^{+},\nu_{e})^{18}O$. For the thermal process, during the accretion phase, the WD has a high density in both its interior and exterior layers, the neutrino production is dominated by plasma decay. Once the TNR occurs, neutrino production will be dominated by photon-neutrino interaction because the temperature of layers is in the temperature range $\rm(T<4\times 10^{8}\ K)$ that it dominates, as shown in the right panel.
@@
Figure 3: The nuclear reactions rates of the nine neutrinos producing nuclear reactions in the pp chain and the CNO cycle during the whole nova multicycle evolution (left). Right: the energy yield of four thermal processes during the whole nova multicycle evolution. The blue line is plasma decay, yellow line is bremsstrahlung, green line is the photon-neutrino interaction, cyan line is electro-pair annihilation. The nova model is similar to Fig. 2.
@@
III.2 Neutrino luminosity of novae with different $M_{\mathrm{WD}}$, accretion rate $\dot{M}$, $T_{\mathrm{C}}$, and mixing
@@
The $M_{\mathrm{WD}}$ of the nova can affect temperatures in the nuclear burning region. The left panel of Fig. 4 shows the neutrino luminosity curves of the range of CO $M_{\mathrm{WD}}$ is $0.6\ \rm M_{\odot}-1.1\ \rm M_{\odot}$. The right panel shows the average peak low-energy nuclear neutrino luminosity of ten outbursts. It is clearly illustrates that the peak luminosity of neutrinos increases as the mass of the WD increases. The larger the mass and smaller the radius of a WD, the greater the gravitational acceleration. This makes it harder for hydrogen to penetrate the degenerate WD core. That means a more degenerate hydrogen-rich envelope, higher temperatures, and, of course, a more violent explosion. As discussed above, the nuclear neutrino is mainly produced by the CNO cycle, which is highly sensitive to template ($\varepsilon_{\rm CNO}\propto T^{17}$). So the more massive the $M_{\mathrm{WD}}$, the higher the low-energy nuclear neutrino luminosity during the TNR.
@@
Figure 4: Neutrino luminosity curves with different CO nova mass models. Left: the neutrino luminosity of the burst sequence during 0--400000 yr with different CO nova mass models, i.e., 0.6 $\rm M_{\odot}$ (cyan), 0.7 $\rm M_{\odot}$ (gray), 0.8 $\rm M_{\odot}$ (red), 0.9 $\rm M_{\odot}$ (green), 1.0 $\rm M_{\odot}$ (blue), and 1.1 $\rm M_{\odot}$ (yellow). Right: average light curves. Both nova models have center temperature $T_{\mathrm{C}}=3\times 10^{7}\ \rm K$, and accrete $25\%$ WD material and $75\%$ solar material at a rate $\dot{M}=1\times 10^{-9}\ \mathrm{M}_{\odot}\rm\ yr^{-1}$.
@@
The left panel of Fig. 5 shows the neutrino luminosity curves of nova models with different $\dot{M}$. The right panel shows the average peak low-energy nuclear neutrino luminosity of ten outbursts. It clearly shows that the higher the accretion rate $\dot{M}$, the smaller the outburst interval $\Delta t_{\mathrm{rec}}$ and $L_{\nu}$. TNR is produced when the hydrogen-rich material is heated to a high enough temperature $\rm(7\times 10^{7}\ K)$ for the hydrogen to burn through the CNO cycle. The high accretion rate provides heat at a high rate [60], thereby inducing an acceleration in nova eruptions. However, owing to the reduced mass of accretion, the lower TNR peak temperature eventually gives rise to a diminished low-energy nuclear neutrino luminosity at higher accretion rates compared to lower ones. When the accretion rate is low, the accretion layer temperature rises slowly and the outburst interval is larger, which means that the accretion layer accretes more material, the outburst intensity is higher, and the low-energy nuclear neutrino luminosity produced during the outburst is higher.
@@
Figure 5: Neutrino luminosity curves with different $\dot{M}$. Left: the neutrino luminosity of the burst sequence during 0--300000 yr with different $\dot{M}$, i.e., $1\times 10^{-9}$ (cyan), $2\times 10^{-9}$ (gray), $3\times 10^{-9}$ (red), $4\times 10^{-9}\ \mathrm{M}_{\odot}\rm\ yr^{-1}$ (green). Right: average light curves. Both nova models have center temperature $T_{\mathrm{C}}=3\times 10^{7}\ \rm K$, and accrete $25\%$ WD material and $75\%$ solar material at a 0.9 $\rm M_{\odot}$ CO WD.
@@
The left panel of Fig. 6 displays the neutrino luminosity curves of a hot model $(T_{\mathrm{C}}=3\times 10^{7})$ and cold models $(T_{\mathrm{C}}=2\times 10^{7}\ \mathrm{and}\ T_{\mathrm{C}}=1\times 10^{7}\ \rm K)$ and with $0.9M_{\odot}$ WD. The right panel shows the average peak low-energy nuclear neutrino luminosity of ten outbursts. When the initial core temperature of the nova model WD is lower, most of the gravitational energy released by the accretion compressed material will be directed to the core, which will delay the time to reach the TNR. Of course, it means the envelope will accrete more H-rich materials, the outbursts are more intense, and the low-energy nuclear neutrino luminosity is higher. However, due to the lower temperature of the core and accreted layer, the nuclear reaction rate during accretion is also lower, resulting in a low neutrino luminosity during the accretion phase. In the hot WD $(T_{\mathrm{C}}=3\times 10^{7}\ \rm K)$, convection zones will appear in the outer core layer. The accreted hydrogen rapidly mixes in the convective region and continues to diffuse to the core, leading to an early outburst [61]. However, because the accretion mass is less and the temperature is lower, the low-energy nuclear neutrino luminosity produced by the outburst is lower.
@@
Figure 6: Neutrino luminosity curves with different $T_{\mathrm{C}}$. Left: the neutrino luminosity of the burst sequence during 0--50000 yr with different $T_{\mathrm{C}}$, i.e., $1\times 10^{7}$ (red), $2\times 10^{7}$ (gray), $3\times 10^{7}\rm\ K$ (cyan). Right: average light curves. Both nova models have accretion rate $1\times 10^{-8}\ \mathrm{M}_{\odot}\rm\ yr^{-1}$, and accrete $25\%$ WD material and $75\%$ solar material at a 0.9 $\rm M_{\odot}$ CO WD.
@@
The left panel of Fig. 7 shows the neutrino luminosity curves of nova models when mixing is considered or not. The right panel shows the average peak low-energy nuclear neutrino luminosity of ten outbursts. If the mixing of accretion material and core CO-enriched material is considered during the accretion stage, it means an increase in the metallicity of the accretion material. On the one hand, increasing the metallicity of the accreting material results in an increase in the opacity. The increased opacity not only allows more heat from the compression of accretion material to be trapped in the accretion layer, but also more heat from nuclear burning to be trapped in the nuclear reaction zone so that the temperature in the nuclear burning region increases faster and the time to reach TNR is significantly shorter than not considering the mixing nova model. Therefore, the total amount of accreted and ejected mass is less [89]. On the other hand, mixing of CO-enriched materials can also outburst earlier because it acts as a catalyst for the CNO cycle (H burning). Models that do not take into account mixing require more accretion matter because the temperature rises more slowly in the nuclear burning region. Thus, compared to nova models that consider mixing, the nova model that does not consider mixing has a denser accretion layer, more intense outbursts, and higher low-energy nuclear neutrino luminosity.
@@
Figure 7: Neutrino luminosity curves with different mixing. Left: the neutrino luminosity of the burst sequence during 0--15000 yr when accreting $25\%$ WD material and $75\%$ solar material and not considering any mixing. Right: average light curves. Both nova models have center temperature $T_{\mathrm{C}}=3\times 10^{7}\ \rm K$, and 0.8 $\rm M_{\odot}$ CO WD at a rate $\dot{M}=1\times 10^{-8}\ \mathrm{M}_{\odot}\rm\ yr^{-1}$.
III.3 Neutrino luminosity of T CrB nova
[…]